# Is Noise Conditioning Necessary for Denoising Generative Models?

### v1

@@ C.1 Proof of the Effective Target
Consider a standard Gaussian variable ${\mathbf{a}}\sim\mathcal{N}(\bm{0},{\bm{I}}_{d})$. Then
@@
$\lim_{d\to\infty}\mathrm{Var}\left[\frac{d}{\|{\mathbf{a}}\|}\right]=\frac{1}{2}.$ (34)
For ease of reading, we will leave the proof of this lemma to the end. Now, we can use the lemma and estimate
$\mathrm{Var}_{t\sim p(t|{\mathbf{z}})}[t]$ $=\mathrm{Var}_{\|{\bm{\epsilon}}^{\prime}\|}\left[\frac{\|{\mathbf{z}}^{\prime}\|}{\|{\bm{\epsilon}}^{\prime}\|}\right]=\left(\frac{\|{\mathbf{z}}^{\prime}\|}{d/2}\right)^{2}\cdot\left(\frac{1}{\sqrt{2}}\right)^{2}\mathrm{Var}_{r:=\frac{\|{\bm{\epsilon}}^{\prime}\|}{\sqrt{2}}}\left[\frac{d/2}{r}\right]$
$\approx\frac{\|{\mathbf{z}}^{\prime}\|^{2}}{d^{2}}\cdot\frac{1}{2}\cdot\frac{1}{2}=\frac{\|{\mathbf{z}}^{\prime}\|^{2}}{4d^{2}}.$ (35)
Finally, we notice that with a high probability, we have
$\|{\mathbf{z}}^{\prime}\|\approx t_{*}\|{\bm{\epsilon}}_{*}^{\prime}\|\approx t_{*}\sqrt{\frac{d}{2}}.$ (36)
Plugging Eq. 36 into Section C.2, we arrive at the desired result Eq. 27. ∎
Proof of Lemma 1. The distribution of the norm of a standard Gaussian variable $r:=\|{\mathbf{a}}\|$ is given by
$p(r)=\frac{1}{2^{\frac{d-2}{2}}\Gamma\left(\frac{d}{2}\right)}r^{d-1}e^{-\frac{r^{2}}{2}},$ (37)
where $\Gamma$ is the gamma function. Then we can directly compute the variance of $d/r$:
$\mathrm{Var}\left[\frac{d}{r}\right]$ $=\int_{0}^{\infty}\left(\frac{d}{r}\right)^{2}p(r)\mathop{}\!\mathrm{d}r-\left(\int_{0}^{\infty}\frac{d}{r}p(r)\mathop{}\!\mathrm{d}r\right)^{2}=\frac{d^{2}}{2}\left(\frac{2}{d-2}-\left(\frac{\Gamma\left(\frac{d-1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}\right)^{2}\right),$ (38)
@@
$\frac{\Gamma\left(\frac{d-1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}$ $=\frac{e^{-\frac{d-1}{2}}\left(\frac{d-1}{2}\right)^{\frac{d}{2}-1}\left(1+\frac{1}{12}\cdot\frac{2}{d}+o(\frac{1}{d})\right)}{e^{-\frac{d}{2}}\left(\frac{d}{2}\right)^{\frac{d-1}{2}}\left(1+\frac{1}{12}\cdot\frac{2}{d}+o\left(\frac{1}{d}\right)\right)}=\sqrt{\frac{2}{d}}\left(1+\frac{3}{4d}+o\left(\frac{1}{d}\right)\right).$ (39)
Plugging Eq. 39 into Eq. 38, we yield
$\lim_{d\to\infty}\mathrm{Var}\left[\frac{d}{r}\right]$ $=\lim_{d\to\infty}\frac{d^{2}}{2}\left(\frac{2}{d}\left(1+\frac{2}{d}+o\left(\frac{1}{d}\right)\right)-\frac{2}{d}\left(1+\frac{3}{2d}+o\left(\frac{1}{d}\right)\right)\right)$
$=\frac{1}{2},$ (40)
finishing the proof. ∎
@@ C.3 Approximation of $E({\mathbf{z}})$
Statement 2  (Error of effective regression targets).  Consider a single datapoint ${\mathbf{x}}\in[-1,1]^{d}$, ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}})$, $t\sim\mathcal{U}[0,1]$, and ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$ (as in Flow Matching). Define $R({\mathbf{z}})$ and $R({\mathbf{z}}|t)$ with the Flow Matching configuration in Table 8. Given a noisy image ${\mathbf{z}}=(1-t_{*}){\mathbf{x}}+t_{*}{\bm{\epsilon}}$ produced by a given $t_{*}$, the mean squared error $E({\mathbf{z}})$ in Eq. 10 can be approximated by
@@
Combining Sections C.3 and 46, we get the estimation in Eq. 41. ∎
@@ C.4 Bound of accumulated error
Statement 3  (Bound of accumulated error).  Starting from the same noise ${\mathbf{x}}_{0}={\mathbf{x}}_{0}^{\prime}$, consider a sampling process (Eq. 4) of $N$ steps, with noise conditioning:
@@
where $\prod_{k=j+1}^{N-1}a_{k}$ is defined as $1$ for $j=N-1$.
@@ Appendix D Derivation of Coefficients for Different Denoising Generative Models
In this section, we build upon our formulation in Section 3.1 to express common diffusion models—iDDPM (Nichol & Dhariwal, 2021), DDIM (Song et al., 2021a), EDM (Karras et al., 2022), Flow Matching (FM) (Lipman et al., 2023; Liu et al., 2023), and our uEDM Model—using a unified notation. The coefficients corresponding to each model are summarized in Table 8 and Table 9, followed by a concise derivation of their formulations.
@@
The loss function of iDDPM (Nichol & Dhariwal, 2021) in DDPM (Ho et al., 2020)’s notation is
@@
Table 8: The coefficients of different models. For iDDPM, we assume a cosine diffusion schedule $\bar{\alpha}(t)$. For both iDDPM and DDIM we follow the original notation of DDPM (Ho et al., 2020). Also note that for EDM, all coefficients are calculated according to first-order ODE solver, and in the final step we need to multiply the output by $\sigma_{\mathop{}\!\mathrm{d}}$ to get the final image. See Appendix D for more details and derivations.
@@
$a(t)$ $\sqrt{\bar{\alpha}(t)}$ $\sqrt{\bar{\alpha}(t)}$ $\frac{1}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $1-t$
@@
$c(t)$ 00 00 $\frac{t}{\sigma_{\mathop{}\!\mathrm{d}}\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $-1$
$d(t)$ $1$ $1$ $-\frac{\sigma_{\mathop{}\!\mathrm{d}}}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $1$
$w(t)$ $1$ $1$ $1$ $1$
@@
Notice the presence of the diffusion schedule $\bar{\alpha}(t)$ inside the coefficients. We adapt a modified version of the cosine schedule in Nichol & Dhariwal (2021):
@@ D.2 DDIM
DDIM (Song et al., 2021a) shares the training process with DDPM (Ho et al., 2020). However, we choose to use the linear schedule for $\bar{\alpha}(t)$, to demonstrate the generality of our scheme. This schedule has the form
@@
Moreover, we can consider the generalized sampler proposed by Song et al. (2021a), which contains an adjustable parameter $\lambda\in[0,1]$. In their original notation, the sampler can be written as
@@ D.3 EDM
The original EDM (Karras et al., 2022) training objective is given by
@@
Finally, the sampling time step is also explicitly given in Karras et al. (2022), so we can directly use it here:
@@ D.4 Flow Matching
The training process of FM (Lipman et al., 2023) is given by
@@ D.5 Our uEDM Model in the Formulation
Introduced in Section 5, the uEDM model designed by us is a modified version of EDM (Karras et al., 2022). The only modification is that we change $c_{\text{in}}(t)$ and $c_{\text{out}}(t)$ by
@@
In Section D.3, we have already derived the general form of the coefficients of EDM with functions $c_{\text{in}}(t)$, $c_{\text{out}}(t)$, $c_{\text{skip}}(t)$ and $\lambda(t)$ in Eqs. 65 and 67. Plugging in the new set of these functions, we can then derive the coefficients of uEDM, as shown in Table 9.
@@
EDM $\dfrac{1}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $\dfrac{t}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $\dfrac{t}{\sigma_{\mathop{}\!\mathrm{d}}\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $-\dfrac{\sigma_{\mathop{}\!\mathrm{d}}}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $1$ $\sqrt{\dfrac{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t_{i}^{2}}{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t_{i+1}^{2}}}\left(1{-}\dfrac{t_{i}(t_{i}{-}t_{i+1})}{t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}\right)$ $\dfrac{\sigma_{\mathop{}\!\mathrm{d}}(t_{i}-t_{i+1})}{\sqrt{(t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2})(t_{i+1}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2})}}$
uEDM $\dfrac{1}{\sqrt{t^{2}+1}}$ $\dfrac{t}{\sqrt{t^{2}+1}}$ $\dfrac{t^{2}}{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}$ $-\dfrac{t\sigma_{\mathop{}\!\mathrm{d}}^{2}}{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}$ $\dfrac{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t^{2}}{\sigma_{\mathop{}\!\mathrm{d}}t}$ $\sqrt{\dfrac{t_{i}^{2}+1}{t_{i+1}^{2}+1}}\left(1{-}\dfrac{t_{i}(t_{i}{-}t_{i+1})}{t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}\right)$ $\dfrac{t_{i}-t_{i+1}}{t_{i}\sqrt{t_{i+1}^{2}+1}}$

### v2

@@ C.1 Proof of the Effective Target
Suppose that ${\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{N}$ are i.i.d Gaussian samples from $\mathcal{N}(\bm{0},{\bm{I}}_{d})$, and ${\mathbf{z}}=(1-t_{0}){\mathbf{x}}_{i}+t{\bm{\epsilon}}$ for some $i$.
Then, we have the following two properties hold with probability $1-\frac{1}{d}$:
$\begin{cases}|\left<{\mathbf{x}}_{j},{\mathbf{z}}\right>|&=\mathcal{O}(\sqrt{d}(\log N+\log d)),\qquad\forall j\neq i\\ |\left<{\mathbf{x}}_{i},{\mathbf{z}}\right>-(1-t_{0})|{\mathbf{x}}_{i}|^{2}|&=\mathcal{O}(\sqrt{d\log d})\end{cases}$
Proof.
We first deal with the second inequality. We have
$\left<{\mathbf{x}}_{i},{\mathbf{z}}\right>-(1-t_{0})|{\mathbf{x}}_{i}|^{2}=\left<{\mathbf{x}}_{i},(1-t_{0}){\mathbf{x}}_{i}+t{\bm{\epsilon}}\right>-(1-t_{0})|{\mathbf{x}}_{i}|^{2}=t_{0}\left<{\mathbf{x}}_{i},{\bm{\epsilon}}\right>\sim\mathcal{N}(\bm{0},|{\mathbf{x}}_{i}|I_{d})$
Using standard Gaussian tail bounds, this probability that $|\left<{\mathbf{x}}_{i},{\mathbf{z}}\right>-(1-t_{0})|{\mathbf{x}}_{i}|^{2}|\geq k\sqrt{d\log d}$ is at most
$\exp\left(-\frac{k^{2}d\log d}{|{\mathbf{x}}_{i}|^{2}}\right)=\exp(-k^{2}\log d)=d^{-\Omega(k^{2})}$
Therefore, there exists a constant $k$ such that the probability above is at most $\dfrac{1}{dN}$.
For the first inequality, notice that
$\left<{\mathbf{x}}_{j},{\mathbf{z}}\right>=(1-t)\left<{\mathbf{x}}_{i},{\mathbf{x}}_{j}\right>+t\left<{\mathbf{x}}_{j},{\bm{\epsilon}}\right$>
each term is the dot product of two normal Gaussian variables, which is still sub-exponential.
In this way, we derive that each term is at most $\mathcal{O}(d(\log N+\log d))$ with probability $1-\dfrac{1}{dN}$, so union bound over all the $i$ gives the desired result. ∎
Theorem 4.
There exists a constant $C$ such that when $1-t_{0}\geq C\sqrt{\dfrac{\log N+\log d}{d}}$, we can recover the correct $i$ with probability at least $1-\dfrac{1}{d}$.
Proof.
Using Lemma 1, we can compute the dot products of all ${\mathbf{x}}_{i}$ with $\mathbf{z}$} and take the largest one. ∎
This theorem implies that the concentration property of $p(t|{\mathbf{z}})$ can be extended to multiple data points case, even when there are exponentially-many data points. By first recovering the correct clean image, the multi–data–point setting reduces immediately to the single–point analysis, and hence the posterior variance bound $\mathcal{O}(t_{0}^{2}/d)$ extends to the case of exponentially many candidates.
C.2.2 Intuitive derivation of the approximate constant
In this section, we show the reason why we estimate the variance as $t_{0}^{2}/2d$.
Firstly, as we have seen in Theorem 4, the ground-truth clean image can be recovered with high confidence, even with a prior of exponentially-many data points. As a result, we revert to the single data point setting.
For simplicity, we restrict attention to the bulk regime $t\in\left[\dfrac{1}{d},1-\dfrac{1}{d}\right]$, postponing edge–case analysis to future work.
Now, we can still view the data point $\mathbf{x}$} as $(x,0,\ldots,0)$ due to symmetry, where $x$ stands for the norm of $\mathbf{x}$}. Because the dominant contribution to the likelihood comes from the d1d-1 noise dimensions when $d$ is large, we further approximate by neglecting the signal component in the first coordinate (i.e., assume ${\mathbf{x}}=\bm{0}$).
On this assumption, we have
${\mathbf{z}}^{\prime}=(1-t){\mathbf{x}}^{\prime}+t{\bm{\epsilon}}^{\prime}=t{\bm{\epsilon}}.$
and denote $|{\mathbf{z}}|=t_{0}\sqrt{d}$. We have
$p(t|{\mathbf{z}})\ \propto\ \frac{1}{t^{d}}\exp\left(-\frac{dt_{0}^{2}}{2t^{2}}\right)$
The variance of this distribution can be calculated exactly from integration, as long as we extend the distribution from $[0,1]$ to the whole $\mathbb{R}$:
@@ C.2.2 Intuitive derivation of the approximate constant
$\mathbb{E}_{p(t|{\mathbf{z}})}[t]=\frac{\displaystyle\int_{-\infty}^{\infty}\frac{1}{t^{d-1}}\exp\left(-\frac{dt_{0}^{2}}{2t^{2}}\right)}{\displaystyle\int_{-\infty}^{\infty}\frac{1}{t^{d}}\exp\left(-\frac{dt_{0}^{2}}{2t^{2}}\right)}=\frac{\displaystyle\frac{1}{2}\left(\frac{2}{dt_{0}^{2}}\right)^{\frac{d-2}{2}}\Gamma\left(\frac{d-2}{2}\right)}{\displaystyle\frac{1}{2}\left(\frac{2}{dt_{0}^{2}}\right)^{\frac{d-1}{2}}\Gamma\left(\frac{d-1}{2}\right)}=\sqrt{\frac{d}{2}}t_{0}\frac{\displaystyle\Gamma\left(\frac{d-2}{2}\right)}{\displaystyle\Gamma\left(\frac{d-1}{2}\right)}$
and similarly
$\mathbb{E}_{p(t|{\mathbf{z}})}[t^{2}]=\frac{dt_{0}^{2}}{2}\frac{\Gamma\left(\frac{d-3}{2}\right)}{\Gamma\left(\frac{d-1}{2}\right)}=dt_{0}^{2}\cdot\frac{1}{d-3}$
which yields
$\text{Var}_{p(t|{\mathbf{z}})}[t]=\frac{dt_{0}^{2}}{2}\cdot\left(\frac{2}{d-3}-\left(\frac{\Gamma(\frac{d-2}{2})}{\Gamma(\frac{d-1}{2})}\right)^{2}\right).$ (32)
Again, here we extend the distribution of $p(t|{\mathbf{z}})$ from $[0,1]$ to $\mathbb{R}$, as this relaxation will only increase the variance.
@@
$\frac{\Gamma\left(\frac{d-2}{2}\right)}{\Gamma\left(\frac{d-1}{2}\right)}$ $=\frac{e^{-\frac{d-2}{2}}\left(\frac{d-2}{2}\right)^{\frac{d-3}{2}}\left(1+\frac{1}{12}\cdot\frac{2}{d}+o(\frac{1}{d})\right)}{e^{-\frac{d-1}{2}}\left(\frac{d-1}{2}\right)^{\frac{d-2}{2}}\left(1+\frac{1}{12}\cdot\frac{2}{d}+o\left(\frac{1}{d}\right)\right)}=\sqrt{\frac{2}{d-1}}\left(1+\frac{3}{4(d-1)}+o\left(\frac{1}{d}\right)\right).$ (33)
Plugging Eq. 33 into Eq. 32, we derive our final estimation:
$\text{Var}_{p(t|{\mathbf{z}})}[t]$ $=\frac{dt_{0}^{2}}{2}\cdot\left(\frac{2}{d-3}-\frac{2}{d-1}\left(1+\frac{3}{2(d-1)}+o\left(\frac{1}{d}\right)\right)\right)$
$=\frac{t_{0}^{2}}{d}\left(\frac{1}{2}+o\left(\frac{1}{d}\right)\right),$
which aligns with our Fig. 3.
@@ C.3 Approximation of $E({\mathbf{z}})$
Statement 1  (Error of effective regression targets). Consider a single datapoint ${\mathbf{x}}\in[-1,1]^{d}$, ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}})$, $t\sim\mathcal{U}[0,1]$, and ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$ (as in Flow Matching). Define $R({\mathbf{z}})$ and $R({\mathbf{z}}|t)$ with the Flow Matching configuration in Table 8. Given a noisy image ${\mathbf{z}}=(1-t_{*}){\mathbf{x}}+t_{*}{\bm{\epsilon}}$ produced by a given $t_{*}$, the mean squared error $E({\mathbf{z}})$ in Eq. 10 can be approximated by
@@
Combining Sections C.3 and 39, we get the estimation in Eq. 34. ∎
@@ C.4 Bound of Accumulated Error
Statement 2  (Bound of accumulated error). Starting from the same noise ${\mathbf{x}}_{0}={\mathbf{x}}_{0}^{\prime}$, consider a sampling process (Eq. 4) of $N$ steps, with noise conditioning:
@@
where $\prod_{k=j+1}^{N-1}a_{k}$ is defined as $1$ for j=N1j=N-1.
@@ Appendix D Derivation of Coefficients for Different Denoising Generative Models
In this section, we build upon our formulation in Section 3.1 to express common diffusion models—iDDPM (nichol2021iddpm), DDIM (song2021ddim), EDM (karras2022edm), Flow Matching (FM) (lipman2023flow; liu2023flow), and our uEDM Model—using a unified notation. The coefficients corresponding to each model are summarized in Table 8 and Table 9, followed by a concise derivation of their formulations.
@@
The loss function of iDDPM (nichol2021iddpm) in DDPM (ho2020denoising)’s notation is
@@
Table 8: The coefficients of different models. For iDDPM, we assume a cosine diffusion schedule $\bar{\alpha}(t)$. For both iDDPM and DDIM we follow the original notation of DDPM (ho2020denoising). Also note that for EDM, all coefficients are calculated according to first-order ODE solver, and in the final step we need to multiply the output by $\sigma_{\mathop{}\!\mathrm{d}}$ to get the final image. See Appendix D for more details and derivations.
@@
$a(t)$ $\sqrt{\bar{\alpha}(t)}$ $\sqrt{\bar{\alpha}(t)}$ $\frac{1}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ 1t1-t
@@
$c(t)$ 0 0 $\frac{t}{\sigma_{\mathop{}\!\mathrm{d}}\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ 1-1
$d(t)$ 11 11 $\frac{\sigma_{\mathop{}\!\mathrm{d}}}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ 11
$w(t)$ 11 11 11 11
@@
Notice the presence of the diffusion schedule $\bar{\alpha}(t)$ inside the coefficients. We adapt a modified version of the cosine schedule in nichol2021iddpm:
@@ D.2 DDIM
DDIM (song2021ddim) shares the training process with DDPM (ho2020denoising). However, we choose to use the linear schedule for $\bar{\alpha}(t)$, to demonstrate the generality of our scheme. This schedule has the form
@@
Moreover, we can consider the generalized sampler proposed by song2021ddim, which contains an adjustable parameter $\lambda\in[0,1]$. In their original notation, the sampler can be written as
@@ D.3 EDM
The original EDM (karras2022edm) training objective is given by
@@
Finally, the sampling time step is also explicitly given in karras2022edm, so we can directly use it here:
@@ D.4 Flow Matching
The training process of FM (lipman2023flow) is given by
@@ D.5 Our uEDM Model in the Formulation
Introduced in Section 5, the uEDM model designed by us is a modified version of EDM (karras2022edm). The only modification is that we change $c_{\text{in}}(t)$ and $c_{\text{out}}(t)$ by
@@
In Section D.3, we have already derived the general form of the coefficients of EDM with functions $c_{\text{in}}(t)$, $c_{\text{out}}(t)$, $c_{\text{skip}}(t)$ and $\lambda(t)$ in Eqs. 58 and 60. Plugging in the new set of these functions, we can then derive the coefficients of uEDM, as shown in Table 9.
@@
EDM $\dfrac{1}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $\dfrac{t}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $\dfrac{t}{\sigma_{\mathop{}\!\mathrm{d}}\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ $\dfrac{\sigma_{\mathop{}\!\mathrm{d}}}{\sqrt{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}}$ 11 $\sqrt{\dfrac{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t_{i}^{2}}{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t_{i+1}^{2}}}\left(1{-}\dfrac{t_{i}(t_{i}{-}t_{i+1})}{t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}\right)$ $\dfrac{\sigma_{\mathop{}\!\mathrm{d}}(t_{i}-t_{i+1})}{\sqrt{(t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2})(t_{i+1}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2})}}$
uEDM $\dfrac{1}{\sqrt{t^{2}+1}}$ $\dfrac{t}{\sqrt{t^{2}+1}}$ $\dfrac{t^{2}}{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}$ $\dfrac{t\sigma_{\mathop{}\!\mathrm{d}}^{2}}{t^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}$ $\dfrac{\sigma_{\mathop{}\!\mathrm{d}}^{2}+t^{2}}{\sigma_{\mathop{}\!\mathrm{d}}t}$ $\sqrt{\dfrac{t_{i}^{2}+1}{t_{i+1}^{2}+1}}\left(1{-}\dfrac{t_{i}(t_{i}{-}t_{i+1})}{t_{i}^{2}+\sigma_{\mathop{}\!\mathrm{d}}^{2}}\right)$ $\dfrac{t_{i}-t_{i+1}}{t_{i}\sqrt{t_{i+1}^{2}+1}}$